OpenStudy (anonymous):

A bag of ten tennis balls that are three different colors. Chaz picks a ball from the bag 40 times without looking and returns it to the bag each time. The results of his experiment are 4 pink, 21 yellow, and 15 purple. How many of each color ball do you think are in the bag? Explain your answers.

OpenStudy (anonymous):

I'd say 5, 4, and 1. This is because the observed probability shows that the chances of picking a pink ball out of the bank are 4/40 = 1/10. Multiplying 1/10 by 10 gives us 1. I multiplied by 10 because that's the total number of balls. Then the observed probability of yellow balls is 21/40. Multiplying by 10 gives us 21/4 which is about 5 since we can't have a fraction of a ball. Then the observed probability of the purple ball is 15/40 = 3/8. Multiplying it by 10 yields 30/8 = 15/4. This is about 4 balls. Hence, there is approximately 1 pink ball, 5 yellow balls, and 4 purple balls.

OpenStudy (anonymous):

Ok Thank you very much